### v1

# Connection between gravity and quantum mechanics: an algebraic approach to cosmological issues

Abstract
This study presents an algebraic framework to explore the fundamental relationship between gravity and quantum mechanics, with particular emphasis on the role of primordial black holes (PBHs) in cosmology. Through the concept of self-gravitating condensed light, specifically in the form of a photon Bose-Einstein condensate, this work examines the quantum attributes of PBHs and their implications for early universe dynamics, baryogenesis, and the formation of large-scale structures. The model also investigates quantized characteristics of black holes, such as mass, entropy, and temperature, suggesting that quantum processes are fundamental to black hole mechanics. By addressing issues like the cosmological constant problem and the information loss paradox, this work provides insights into Planck-scale physics and proposes that PBHs may serve as a bridge between quantum theory and general relativity. This study ultimately posits that understanding black hole physics is essential to resolving major cosmological and astrophysical paradoxes for the ultimate unification of quantum mechanics with gravity.
I Introduction
This paper aims to describe how primordial black holes are significant in the formation and evolution of the observable universe. Besides, it is trying to solve and interpret numerous questions and paradoxes relevant to Planck-scale physics, Big Bang cosmology, baryogenesis, the mechanism of Supermassive black holes (SMBHs), the early clumpiness of the universe, large-scale structures, such as supervoids and superclusters, and two-stage accelerated expansion; the horizon problem, cosmological constant problem, and the information-loss paradox that has arisen in observational cosmology and astrophysics. We shall demonstrate that the realistic physics of black holes is relevant to understanding fundamental cosmology problems. To clarify the above issues, let us first introduce the connection between the physics of primordial black holes and Bose-Einstein condensates at the Planck scale limit. This will be the main subject of the paper.
The fundamental discovery of a previously unknown state of matter, self-gravitating condensed light [1, 2], played an outstanding role in our research. We are discussing the quantized photon Bose-Einstein condensate [2, 3, 4], a phenomenon predicted approximately a century ago. Let us discuss the quantum nature of self-gravitating condensed light.
Within a vacuum, photons fail to be involved in direct interaction according to conventional quantum electrodynamics (QED). Photons often behave as though they interact via effective mass or higher-order quantum corrections (e.g., via virtual particle-antiparticle pairs in quantum loops). However, this may change in extreme situations like nonlinear media or close to strong gravitational fields. A self-gravitating system is possible because of these interactions, which allow photons to produce their gravitational fields potentially. The Euler-Heisenberg Lagrangian [1], which characterizes non-linear interactions at high energy scale QED, is the source of the idea of photon-photon scattering. Photons in such regimes may change the spacetime curvature around them and behave as gravity sources via self-interactions, which significantly change the behavior of light.
Bose-Einstein condensation (BEC), primarily associated with massive bosons such as atoms, can also be applicable to photons under specific conditions. Photons, as bosons, are capable of occupying identical quantum states, resulting in a macroscopic quantum phenomenon characterized by the presence of numerous photons in a coherent, condensed state. Considering the influence of gravity, these condensed photons may have the capacity to establish a self-gravitating system. Photons confined in optical cavities or thermalized through interactions with matter can attain BEC at temperatures above absolute zero [2]. At high energy densities, a self-bound photon configuration may theoretically emerge, similar to the formation of gravitationally bound systems such as stars or planets from massive particles.
The self-gravitating condensed light hypothesis introduces the idea of photon stars, which are gravitationally bound systems composed entirely of photons. Such theoretical objects are also known as geons, as they were conceived as solutions to Einstein’s field equations. Concentrated electromagnetic radiation would induce enough gravity to prevent the dispersion of photons [5]. Such stars are stable because the gravitational attraction is balanced by quantum pressure from photons in a Heisenberg uncertainty relation analogous to that for neutron stars, stabilizing the latter against collapse by neutron degeneracy pressure.
There is a balance between gravitational forces and quantum mechanical effects, especially the Heisenberg uncertainty principle, that makes a condensed light system stable in self-gravitating quantum frameworks. Quantum pressure, arising from the uncertainty in photon momentum and position, prevents gravitational collapse, similar to degeneracy pressure in white dwarfs and neutron stars. In semiclassical gravity [6], where quantum fields interact with classical spacetime, self-gravitating photon condensates may become stable in some situations. This gives us clues about where quantum field theory and general relativity meets. It might help us learn more about quantum gravity, cosmology, and high-energy astrophysics by showing us how photons behave in strong gravitational fields and how quantum effects happen in spacetime that is curved [7].
The early universe might have existed as a dense, “grainy” spacetime, possibly formed from a Planck-scale photon condensate, as proposed by quantum gravity theories [8, 9]. At this scale, spacetime is believed to be discrete rather than continuous, a notion supported by approaches like Loop Quantum Gravity(LQG) and string theory. These “grainy” structures could reflect quantum fluctuations, suggesting that photons formed a BEC under extreme energy conditions [10]. This condensate, where photons share a collective quantum state, may have acted as the foundational structure of spacetime. The Planck photon 3D condensate concept posits that at extremely high energies, especially near the Planck scale, photons may aggregate to form a BEC [8]. According to the holographic principle, which holds that black hole surfaces encode higher-dimensional information, PBHs can also be considered 2D photon condensates or ”membranes” that exhibit quantum characteristics [11].
The quantum characteristics of self-gravitating condensed light inherently lead to the quantization of the geometrical, gravitational, thermodynamical, informational, and radiational properties of black holes, which are described as follows:
The geometry of black holes is linked to the discrete characteristics of spacetime at the Planck scale. In quantum gravity theories such as LQG, spacetime is characterized as discrete, consisting of distinct “quanta” of space. This framework results in the inherent quantization of geometrical properties [12], including area and volume. In the context of black holes, this indicates that the surface area of the event horizon is quantized in units of the Planck area.
The gravitational field, as articulated by the Einstein field equations, is affected by quantum corrections at small scales [13]. In models of self-gravitating light, quantum fluctuations affect the gravitational field made by condensed photons. This creates a quantized spectrum of gravitational effects, especially in very dense objects like black holes.
Black hole thermodynamics, as articulated by Bekenstein-Hawking entropy [14, 15], demonstrates that black holes possess discrete energy levels, similar to those found in atomic systems. The entropy of a black hole, which is proportional to the area of its event horizon, is quantized in units of the Planck area. This is derived from the quantum mechanical interpretation of black hole horizons, in which microstates represent various quantum configurations of the horizon. Additionally, it should be noted that in his early works, S. Hawking et al. [16, 17] provided particular focus to the exceptional role of PBHs in his cosmological research.
The quantum features associated with black holes result in the quantization of information storage and radiation [18], especially as demonstrated by Hawking radiation. Black holes emit radiation in discrete quanta, and the information capacity of a black hole is proportional to the surface area of its event horizon. The quantization of information is directly connected to the black hole information paradox, wherein the loss of information through Hawking radiation presents an unresolved issue in quantum gravity. The quantized characteristics of black holes illustrate the profound relationship between quantum mechanics and gravity, particularly in extreme conditions where both theories are anticipated to merge into a unified framework.
On the other hand, V. Mukhanov and J. D. Bekenstein developed their research on quantized black holes, enhancing the comprehension of quantum phenomena in black hole physics [19, 20, 21]. Mukhanov’s research concentrated on quantum fluctuations during inflation [22], which subsequently proved essential to the hypothesis of cosmic structure development. His investigation on quantized black holes fostered discourse on the quantum characteristics of black hole horizons and the possibility of PBHs providing an understanding of quantum gravity. It should also be noted that recently, a growing number of astrophysicists contend that dark matter is comprised of microscopic PBHs with asteroid-like masses. A comprehensive review is available in Refs. [23, 24].
Recent remarkable discoveries, including Einstein’s gravitational waves [25, 26, 27, 28], “impossible” early galaxies and quasars [29, 30, 31, 32], and the 2D photon Bose-Einstein condensate with rest energy [2, 3, 35, 36], may offer a strong experimental and observational foundation for what John Wheeler famously described as the “fiery marriage of general relativity with quantum theory.” Let us discuss “impossible” early galaxies and quasars, and the 2D photon Bose-Einstein condensate with rest energy in brief.
Extremely large galaxies and quasars found in the early universe can pose serious questions to present-day models of cosmology. Observations of galaxies and quasars by advanced telescopes, like JWST, show that they have high masses, luminosities, and metallicities that go against current theories about how galaxies form and evolve. These theories are mostly about redshifts greater than $10$ [29, 30]. These so-called “impossible early galaxies” formed less than $500$ million years after Big Bang, indicating star formation rates and galaxy-assembly processes that exceed the predictions of conventional models. This indicates rapid star formation and galaxy assembly beyond the expectations of the standard Lambda Cold Dark Matter ($\Lambda$CDM) model, which actually predicts a gradual build-up of structure. Some galaxies appear massively, to the tune of $10^{9}$ solar masses within $300$ million years, and indicate accelerated star formation or unknown mechanisms [31]. Early quasars that are powered by SMBHs of mass $>10^{9}$ solar masses at redshift $>6$ also pose a challenge, since such black holes should take longer to form at the limits of accretion rates. Recent discoveries of quasars at redshifts $>7$ [32], challenge existing limits, and suggest the presence of “black hole seeds” that have significantly larger mass than those of consideration in earlier models or alternative growth mechanisms, including direct collapse scenarios. These results therefore demand reconsideration of $\Lambda$CDM or alternative theories, such as PBHs, non-standard dark matter interactions, or novel physics in the early universe.
Additionally, Belgiorno et al. [33] describe the analogous Hawking radiation that includes Bose-Einstein condensates and surface gravity waves. The examination of near-horizon phenomena in more controlled laboratory settings can be achieved through these systems that simulate analogous Hawking radiation. This facilitates investigations into grey-body factors, thermality, and dispersive effects, contributing to the experimental validation of Hawking radiation. The recent study conducted by Agullo et al. [34] examines the generation of entangled particle pairs within the context of 2D Bose-Einstein condensate analogs in expanding cosmological models. They have pointed out the significance of entanglement as a critical quantum indicator of pair creation and showcased experimental advancements in the potential to simulate quantum phenomena pertinent to cosmological contexts. These results converge to explain the influences of quantum gravity and its implications for the dynamics of the early universe and the formation of large-scale structures.
On the other hand, the concept of a two-dimensional (2D) photon Bose-Einstein condensate introduces a novel approach to studying the quantum statistical behavior of light by combining quantum optics, condensed matter physics, and thermodynamics. Unlike traditional atomic BECs, where atoms condense at low temperatures, photons—typically massless—can achieve a condensed state by simulating an effective mass within a planar optical microcavity. This is achieved by confining photons between curved mirrors, creating a 2D environment where they behave similarly to massive particles. A heat reservoir, often via dye molecules, allows the photons to reach thermal equilibrium, forming a ”rest energy” and enabling condensation at reduced temperatures [2, 35]. As the system cools, photons transition to the lowest energy mode, forming a coherent quantum state [36]. This 2D photon BEC platform offers significant insights into fundamental physics, potentially impacting low-threshold laser technology, quantum information, and the study of thermodynamic properties in non-equilibrium quantum systems [35]. Recently, Mazur and Mottola [37] introduced the concept of Gravitational Condensate Stars as a stable alternative to black holes. In their work, they present a compact, non-singular solution to Einstein’s equations, suggesting it as a possible endpoint for gravitational collapse. Unlike black holes, this type of collapsed star aligns with quantum theory, remains thermodynamically stable, and avoids the information paradox.
Integrating general relativity and quantum mechanics is crucial to deepening our understanding of the universe and enhancing space exploration. While effective in their domains, these theories must unify to fully explain extreme cosmic phenomena like black holes and the early universe. Gravitational waves (spacetime ripples) and quantum probability waves are fundamentally different yet share a geometric link, suggesting a pathway for unification. This approach involves analyzing phenomena, especially black holes, through wavelengths and geodesics, bridging the two frameworks. A key objective is to identify a quantum “background substance” that balances black holes’ intense gravitational forces with quantum repulsion, shedding light on effects like Hawking radiation and quantum tunneling. This could ultimately deepen our grasp of the universe across both large and small scales.
This manuscript is structured in the following manner: In section II, we derived the algebra (namely Einstein’s algebra) utilizing the membrane formalism, establishing a correspondence between the primordial black holes and the two-dimensional condensate photons. In section III, we examine the essential foundational principles of physics at the Planck scale. In section IV, we elucidate the interconnections between energy, temperature, and entropy of black holes when they approach the Planck scale limit. The discussion of primordial black holes’ lifetime and luminosity, in conjunction with the information loss paradox, is presented in Section V. Subsequently, we outlined the process of free light accretion via quantum absorption, as well as the origins of disks and jets in section VI. In section VII, we analyze the conservation law of quantum information and the energies associated with gravitational radiation from colliding and merging black holes, which serve as a driving force behind the accelerated expansion of the universe. In section VIII, we examine the intricate details of Planck-scale phenomena, the dynamics of the hot Big Bang, and the initial phase of the universe’s accelerated expansion. In section IX, we derived the connection between the Planckian Era and the Baryogenesis Epoch, addressing the issue of supermassive black hole formation and the primordial clumping of the universe. In section X, we analyze baryogenesis via the mechanism of primordial black hole production. In section XI, we established the methodology for the formation of contemporary dark matter and the large-scale structure of the universe. The final section, XII, serves as the discussion and conclusion of this study.
II The sketches of “Einstein’s algebra”
In this section, we begin with Albert Einstein’s final work in 1956 [38], where he articulated a concluding statement that reflects his core methodological principle: “… a finite system of finite energy can be completely described by a finite set of quantum numbers. This does not seem to align with a continuum theory and must lead to an attempt to find a purely algebraic theory for describing reality. However, nobody knows how to establish the foundation for such a theory.”
In 1985, Price and Thorne introduced the ”membrane formalism” for black holes [39], which reframes the black hole’s horizon as a ”stretched horizon”—a 2D membrane within 3D space that responds dynamically to external forces. This membrane, building on Damour and Znajek’s work [40, 41, 42], acts like a viscous fluid with electrical charge and conductivity, as well as finite entropy and temperature, though it cannot conduct heat. The membrane’s interactions with the universe are governed by familiar fluid dynamics and thermodynamic principles, such as the Navier-Stokes equation, Ohm’s law, and the laws of thermodynamics. These concepts provide astrophysicists with intuitive tools for understanding black hole behavior in complex environments. In this context, we discuss the following two works. The discussion covers two works focused on early universe phenomena. In Grasso et al. [43], the authors explore the possibility that galactic and cluster magnetic fields have a primordial origin. They suggest that searching for imprints of these magnetic fields in the temperature and polarization anisotropies of the cosmic microwave background radiation (CMBR) is the best approach for testing this hypothesis. Another work done by Bruce et al. [44] proposes that dark energy could originate from dark matter condensation at a low redshift. They analyze constraints from CMBR, large-scale structure, supernovae, and radio galaxies using a four-parameter likelihood analysis, factoring in the Hubble constant and three parameters of the condensate field, $Q$.
The previously mentioned concepts may also apply to 2D photon Bose-Einstein condensates (BEC) with rest energy in various phenomena [2, 3, 45, 46]. The idea of a 2D photon BEC with rest energy can be linked to the early universe, particularly concerning PBHs, analog gravity, and related topics [47, 48, 49, 50]. This correlation can be framed within a physical model where PBHs are represented as 2D spherical photon condensates contained within their own gravitational fields. Such a model naturally leads to a geometrical equation that directly connects the Compton wavelength, $\lambda_{sm}$, of condensed light (quantum theory) with the geodesic length (general relativity) [4]:
$\lambda_{sm}=2\pi R_{s},$ (1)
where $R_{s}$ is Schwarzschild radius. Based on the above discussions, we can intuitively say that spacetime instructs condensed light quanta on how to curve, while condensed light quanta instructs spacetime on how to quantize. According to the definition of the Compton wavelength and the expression for Schwarzschild radius, we may express it as
$\lambda_{sm}=$ $\frac{2\pi\hbar}{M_{sm}c}$
$2\pi R_{s}=$ $\frac{4\pi GM_{s}}{c^{2}}$ (2)
Using the above relationship Eqs. (1) and (2), we promptly determine the mass and rest energy of each condensed photon:
$M_{sm}=\frac{2\pi\hbar}{\lambda_{sm}c}=\frac{\hbar}{R_{s}c}$
$E_{sm}=M_{sm}c^{2}=\frac{\hbar c}{R_{s}}$ (3)
along with their aggregate quantity:
$N_{s}=\frac{M_{s}}{M_{sm}}=2\frac{M_{s}^{2}G}{\hbar c}=2\frac{M_{s}^{2}}{M_{p}^{2}}=2\frac{E_{s}^{2}}{E_{p}^{2}}=\frac{R_{s}^{2}}{2l_{p}^{2}},\leavevmode\nobreak\ \text{etc.},$. , (4)
expressed in terms of the Planck mass $M_{p}\leavevmode\nobreak\ (=\sqrt{\frac{\hbar c}{G}})$, Planck energy $E_{p}$, and Planck length $l_{p}\leavevmode\nobreak\ (=\sqrt{\frac{\hbar G}{c^{3}}})$. The fundamental and continuous relationships may enable us to explore the unknown facets of Planck-scale physics. Considering that the quantity $N_{s}$ of condensed photons cannot be less than two, we readily derive a significant expression for $N_{s}$, utilizing the natural quantum number $n_{s}$ as
$N_{s}=2n_{s},\leavevmode\nobreak\ n_{s}=1,2,3,...$ (5)
Relying solely on established principles of physics, one can readily determine the quantized characteristics of PBHs - such as mass $M_{s}$, rest energy $E_{s}$, size $R_{s}$, entropy $S_{s}$, temperature $T_{s}$, quantum information quantity $I_{s}$, longevity $\tau_{s}$, luminosity $L_{s}$, among others—as well as the fundamental principles governing their formation, evolution, and eventual decay. It is important to note that all these quantized features of PBHs depend exclusively on Planck units and are influenced by various powers of natural quantum numbers [4]. From the previously established quantum relations we get,
$M_{s}=M_{p}\sqrt{n_{s}};\leavevmode\nobreak\ E_{s}=E_{p}\sqrt{n_{s}};\leavevmode\nobreak\ R_{s}=2l_{p}\sqrt{n_{s}};\leavevmode\nobreak$
$A_{s}=4\pi R_{s}^{2}=16\pi l_{p}^{2}n_{s};\leavevmode\nobreak\ S_{s}=2k_{B}n_{s};\leavevmode\nobreak\ T_{s}=\frac{T_{p}}{4\sqrt{n_{s}}};\leavevmode\nobreak$
$I_{s}=N_{s}=2n_{s}(qubit);\leavevmode\nobreak\ \tau_{s}=\frac{4\pi t_{p}n_{s}^{\frac{3}{2}}}{3};\leavevmode\nobreak\ L_{s}=\frac{L_{p}}{4\pi n_{s}},$ (6)
where $t_{p}$, $L_{p}$, and $T_{p}$ represent the Planckian units of the lifetime, luminosity, and temperature, respectively, and $k_{B}$ denotes the Boltzmann constant. We can easily find expressions for the above parameters by putting $n_{s}=1$ for a minimal Planck black hole.
In this discussion, we focus on the quantum mechanical principles underlying the phenomenon of a black hole (PBHs). By rephrasing $n_{s}$ in Planck scale units, we can enable a direct comparison of our results with the findings drawn by Bekenstein [51] and Hawking [14]:
$S_{s}=\frac{k_{B}c^{3}}{8\pi\hbar G}A_{s}=\frac{S_{B}}{ln2}=\frac{S_{H}}{2\pi}$
$T_{s}=\frac{\hbar c^{3}}{4k_{B}GM_{s}}=2\pi T_{H}$ (7)
where we define $S_{B}\leavevmode\nobreak\ (=\frac{k_{B}c^{3}ln2}{8\pi G\hbar}A)$ to represent Bekenstein entropy [51], $S_{H}\leavevmode\nobreak\ (=\frac{k_{B}c^{3}}{4G\hbar}A)$ to signify Hawking entropy [14], and $T_{H}\leavevmode\nobreak\ (=\frac{\hbar c^{3}}{8\pi GMk_{B}})$ to denote Hawking temperature.
Our findings show that our outcomes closely match those that Bekenstein and Hawking predicted, indicating that we are moving in the right direction. However, our findings were derived not from intuitive assumptions, but through clear and exact physical principles, relationships, and an adequate experimental basis.
III Planck equivalence and corresponding relations
It is well-recognized that a hallmark of any well-established theory is its natural derivation of fundamental constants, equivalences, and relationships. In our work, we independently developed a significant system of Planck units, necessitating a thorough examination of the physical implications and extensive consequences that arise from it. For instance, this approach reveals a series of astonishing formulas and equivalences demonstrating the system’s inherent depth and coherence. The formulas and equivalences are:
[…]
$=\hbar\omega_{p}\leavevmode\nobreak\ (Planck),\leavevmode\nobreak\ \omega_{p}=\frac{1}{t_{p}};$
[…]
$=\frac{\hbar}{l_{p}}\leavevmode\nobreak\ (de\leavevmode\nobreak\ Broglie);$
[…]
where $a_{p}$ denotes Planck acceleration and $\omega_{p}$ signifies Planck cyclic frequency, $g_{p}$ is Planck gravity, and so on. The above relations can be utilized in the subsequent discussion, particularly in the context of the initial state of the observable universe and the processes related to the Big Bang.
IV Energy and entropy radiation of black holes
The description of a black hole closely aligns with the renowned “membrane formalism” proposed by Price and Thorne [39], which reconceptualizes the dynamics of the black hole horizon as a two-dimensional “stretched horizon” that operates as a physical membrane within three-dimensional space. In our case, we identify two significant differences from that perspective [4].
1. We introduced the concept of “membrane stretched horizon” as a two-dimensional photonic condensate.
2. In our framework, the significant gravitational squeezing force, which explained as below:
$F_{s}\leavevmode\nobreak\ (gravitational\leavevmode\nobreak\ squeezing\leavevmode\nobreak\ force)$
$=\frac{dE_{s}}{dR_{s}}=\frac{d}{dR_{s}}(\frac{M_{p}c^{2}R_{s}}{2l_{p}})=\frac{c^{4}}{2G}=\frac{F_{p}}{2}$ (9)
which is precisely balanced by analogous repulsive quantum-mechanical counterpart as:
$F_{s}\leavevmode\nobreak\ (quantum\leavevmode\nobreak\ mechanical\leavevmode\nobreak\ force)$
$=N_{s}\frac{dE_{sm}}{dR_{s}}=\frac{R_{s}^{2}}{2l_{p}^{2}}\frac{d(\frac{\hbar c}{R_{s}})}{dR_{s}}=-\frac{F_{p}}{2}.$ (10)
This has profound implications for forming a stable quantum gravitational state that prevents singularities. An equilibrium between gravitational compression and quantum repulsion can establish conditions that preserve quantum information at the horizon, consistent with concepts from holography and the AdS/CFT correspondence [52, 7, 53, 54]. The Holographic principle, proposed by Gerard ’t Hooft and further developed by Leonard Susskind, proposes that all information contained within a volume can be encoded on its boundary, analogous to a three-dimensional hologram represented on a two-dimensional surface. The AdS/CFT correspondence, introduced by Juan Maldacena in 1999, exemplifies this principle’s application. This work shows how two different theories of gravity work together in AdS/CFT at its boundary. It also makes a link between quantum field theory and gravity, which makes it easier to study quantum gravity, black holes, and condensed matter physics.
However, this remarkable equilibrium is periodically interrupted by spontaneous quantum transitions, as described by the work of Bekenstein and Mukhanov [21]. Let’s delve deeper to understand what this means in the context of the structure formation of the universe. According to our understanding, quantum fluctuations are critical in establishing the initial density variations in the early universe that eventually grow into galaxies and large-scale structures. These fluctuations create a “spontaneous transition” in the nascent universe by introducing new phases or areas with slightly differing energy densities. A “spontaneous quantum transition” could occur when the universe’s energy density reaches a critical threshold, causing it to “jump” to a new state. This could be analogous to vacuum decay or phase transitions in early cosmology, in which spacetime “selects” regions with lower energy, resulting in the expansion or reconfiguration of spacetime domains. These transitions occur from an initial $n_{s}$ energy level to a lower $n_{s}-1$ state, resembling electron transitions in the hydrogen atom. The conservation of momentum drives such transitions, which involve the emission of two particles. Gravitational radiation ($\Gamma_{s}$), which appears in quadrupolar and octupolar wave modes, also coincides with these quantum leaps ($\gamma_{s}$). The fundamental framework governing these transitions is outlined as follows [4]:
$BH_{s}\to BH_{s-1}+2\gamma_{s}+\Gamma_{s}$ (11)
with
$E(BH_{s})=E_{p}\sqrt{n_{s}};\leavevmode\nobreak\ E(BH_{s-1})=E_{p}\sqrt{n_{s}-1}.$ (12)
So that we can write
$\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)=E_{p}\sqrt{n_{s}}-E_{p}\sqrt{n_{s}-1}$ (13)
which can also be written as
$\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)=\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)+\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)$
where $qu(rad)$ and $gr(rad)$ represents quantum radiation and gravitational radiation respectively. Also, note that throughout the article, ($qu$), ($gr$), and ($rad$) represent “quantum”, “gravitational”, and “radiation” respectively. Let us evaluate the following terms:
$\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)$
$=2\times\frac{E_{p}}{4\sqrt{n_{s}}}\leavevmode\nobreak\ (two-particle\leavevmode\nobreak\ energy)$ (15)
and from Eq. (14), we get
$\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)=\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)-\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)$
$=E_{p}\sqrt{n_{s}}-E_{p}\sqrt{n_{s}-1}-\frac{E_{p}}{2\sqrt{n_{s}}}$
$=\Delta E_{s,s-1}\leavevmode\nobreak\ (qu)\big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)}$ (16)
This transition resembles those observed in a hydrogen atom, occurring with the emission of two particles to conserve momentum and resulting in gravitational radiation. Gravitational waves arise from the strong gravitational fields of black holes, leading to the distortion of the surrounding spacetime. The most intense gravitational radiation generated in this process exhibits quadrupolar and octopolar features. The coefficients in Eq. (16), $\frac{1}{4n_{s}}$ and $\frac{1}{8n_{s}^{2}}$, correspond to the quadrupolar and octupolar modes, respectively. From the Clausius relation, we obtain the radiated quantum and gravitational entropy:
$\Delta S_{s,s-1}=\frac{\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)}{T_{s}\leavevmode\nobreak\ (rad)}$
$=\Delta S_{s,s-1}\leavevmode\nobreak\ qu(rad)+\Delta S_{s,s-1}\leavevmode\nobreak\ gr(rad)$ (17)
with
$\Delta S_{s,s-1}\leavevmode\nobreak\ qu(rad)=k_{s}2\leavevmode\nobreak\ (qubit);$
$\Delta S_{s,s-1}\leavevmode\nobreak\ gr(rad)=k_{s}2\leavevmode\nobreak\ (qubit)\big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)},$
where $\Delta E_{s,s-1}\leavevmode\nobreak\ (rad)$ all emitted energy; $\Delta E_{s,s-1}\leavevmode\nobreak\ qu(rad)$ the energy of two emitted particles; $\Delta E_{s,s-1}\leavevmode\nobreak\ gr(rad)$ the energy of gravitational radiation; $T_{s}\leavevmode\nobreak\ (rad)=T_{s}$; $\Delta S_{s,s-1}\leavevmode\nobreak\ (rad)$ the emitted entropy, that increases the total entropy of the universe; $\Delta S_{s,s-1}\leavevmode\nobreak\ qu(rad)$ two particle radiation entropy, $\Delta S_{s,s-1}\leavevmode\nobreak\ gr(rad)$ gravitational radiation entropy. We are typically dealing with the emission of two photons, however, there is one very significant exception to this rule when it comes to baryogenesis (see below).
In order to clarify the profound physical implications of the results obtained, it is essential to “dress” the “naked” mathematical expression $\Delta E_{s,s-1}(rad)$ in explicit physical terms.
The careful testimony towards Planck equivalences and correlations is helpful to us in the following ways:
  1. 1.
    Considering the straightforward physical expression of the acceleration $a_{s}$, linked to the dynamical process of a black hole’s evaporation
    $a_{s}=\frac{F_{s}}{M_{s}}=\frac{\frac{F_{p}}{2}}{M_{p}\sqrt{n_{s}}}=\frac{a_{p}}{2\sqrt{n_{s}}},$ (19)
    one can easily find the very interesting physical form that links emitted energy and acceleration, as mentioned above with the Planck equivalence $E_{p}=\frac{\hbar}{c}a_{p}$:
    $\Delta E_{s,s-1}(rad)$
    $=\frac{\hbar}{c}a_{s}+\frac{\hbar}{c}a_{s}\big{(}\frac{1}{4n_{s}}+\frac{1}{8n_{s}^{2}}+...\big{)}$ (20)
    where the first term expresses the energy of two emitted particles, the second term is related to the energy of gravitational radiation, expressed by quadrupolar $\frac{1}{4n_{s}}$ and octupolar $\frac{1}{8n_{s}^{2}}$ terms.
  2. 2.
    Using the previously stated equivalence $T_{p}=\frac{a_{p}}{ck_{B}}\hbar$, we can derive the remarkable expression that relates the temperature and acceleration as:
    $T_{s}(rad)=\frac{E_{p}}{4ck_{B}\sqrt{n_{s}}}$
    $=\frac{\hbar a_{p}}{4ck_{B}\sqrt{n_{s}}}=\frac{\hbar}{2ck_{B}}a_{s}=T_{s}.$ (21)
    This outcome indicates that the temperature of the quantum state is dependent upon acceleration, complementing Unruh’s finding [55] that an accelerated observer in a vacuum perceives it as a warm bath of particles. For a black hole, the term $\sqrt{n_{s}}$ signifies that the actual temperature varies with the quantum state $n_{s}$, implying a quantized nature to this thermal phenomenon. This connection to Unruh’s theorem underscores the profound interconnection of quantum mechanics, thermodynamics, and relativity, illustrating how the black holes accelerated evaporation influences the observed properties of the vacuum.
It should be noted that the Unruh theorem [55] played a pivotal role in theoretical physics by enabling T. Jacobson [56] to derive Einstein’s field equations using thermodynamic principles. Building on this, Jacobson’s profound idea, often summarized as “black hole entropy without black holes”—led to the application of the Clausius relation in this context, connecting black hole thermodynamics with the broader framework of spacetime geometry. The values of gravitational radiation energy and radiated gravitational entropy are substantial only in the Planck scale range, namely during the Planckian era where quantum numbers commence with the initial natural numbers. However, it is these initial numbers that will unlock the gateway to the Beginning of Time. This has a significant implication in quantum cosmology according to which the universe’s birth involved quantum gravitational phenomena. The quantization of time, space, and gravitational waves at the Planck epoch may imply that classical spacetime evolved from a fundamentally discrete quantum structure. Indeed this idea provides us a way to uncover the quantum structure underlying the universe’s beginning and to gain a deeper understanding of time itself.
V PBHs Lifetime and Luminosity: Addressing the Information Loss Paradox
PBHs provide a distinctive platform to test theories of quantum gravity and black hole evaporation. Here, we aim to establish a connection between general relativity and quantum physics by examining lifetimes, luminosities, and the potential elimination of information-loss paradoxes. The time of transition of PBHs from one quantum state to the underlying one can be calculated using the strong conditions of causality and spherical geometry. It is equal to half the length of the geodesic line divided by the speed of light [4],
$\Delta t_{s}=\frac{\pi R_{s}}{c}=2\pi t_{p}\sqrt{n_{s}}.$ (22)
The aforementioned expression for black hole lifetime can be readily found using this formula as:
$\tau_{s}=\sum_{n=1}^{n_{s}}2\pi t_{p}\sqrt{n}$
[…]
$=\frac{4\pi G^{2}M_{s}^{3}}{3\hbar c^{4}}$
$=\frac{\tau_{s}\leavevmode\nobreak\ (Hawking)}{3840}$ (23)
and Luminosity
$L_{s}=\frac{\Delta E_{s,s-1}}{\Delta t_{s,s-1}}=\frac{\frac{E_{p}}{2\sqrt{n_{s}}}}{2\pi t_{p}\sqrt{n_{s}}}$
$=\frac{L_{p}}{4\pi n_{s}}=\frac{\hbar c^{6}}{4\pi G^{2}M_{s}^{2}}$
$=3840L_{s}\leavevmode\nobreak\ (Hawking),$ (24)
where the usual lifetime and luminosity of the black hole can be expressed as: $\tau(Hawking)\simeq\frac{5120\pi G^{2}}{\hbar c^{4}}M^{3}$ and $L(Hawking)\simeq\frac{\hbar c^{6}}{15360\pi G^{2}M^{2}}$ [14, 15, 57]. This model suggests that black holes undergoing quantum transitions differ substantially from Hawking’s semi-classical evaporation model. Quantum processes result in more intense emissions than continuous black-body radiation, reducing the black hole’s lifetime by a factor of 3840 and increasing its luminosity by the same factor due to discrete energy jumps. Consequently, without external energy input (e.g., through accretion) to offset this intense quantum emission, the black hole undergoes rapid evaporation. This finding challenges Hawking’s assumption that black holes radiate as black bodies under the Stefan-Boltzmann law, showing instead that quantum transitions produce significantly higher radiation intensities.
Now we calculate the energy of each emitted particle:
$E_{\gamma}(rad)=\frac{\Delta E_{s,s-1}qu(rad)}{2}=\frac{E_{p}}{4\sqrt{n_{s}}}.$ (25)
From the above Eq. (25), one can easily find that their emitted wavelength is much more than $R_{s}$ :
$\lambda_{\gamma}(rad)=\frac{2\pi\hbar c}{E_{\gamma}(rad)}=4\pi R_{s}>R_{s}.$ (26)
Because the wavelength of the emitted particles is greater than the Schwarzschild radius, the black hole’s strong gravitational field does not prevent the particles from tunneling out. In our scenario, each quantum transition results in the emission of two particles, each containing a discrete amount of energy and information. As a result, each emission causes the black hole to lose two bits of quantum information. However, in conventional black hole evaporation, Hawking radiation is merely thermal and contains no information about the black hole’s internal state, resulting in the so-called information loss paradox. In contrast, our model may encode information in the emitted particles, allowing information about the black hole’s interior states to escape with each two-particle emission. As the black hole emits radiation, it loses mass, energy, and information in a quantised manner, resolving the information loss dilemma by allowing information to leave the black hole in the form of emitted quantum bits (qubits).
VI Quantum Absorption of Free Light: Origins of Cosmic Disks and Jets
Considering the quantum characteristics of condensed light in black holes, it is evident that only free light photons with a wavelength not exceeding the geodesic length of the black hole are subject to accretion by quantum absorption. In other words, black holes consume only free light or condensed light (such as other black holes through collisions and mergers). All ordinary matter is melted in accretion disks and expelled in jets, accompanied by swirling magnetic fields through “funnel” and “slingshot” effects. In this scenario, the wavelength of an absorbed photon cannot exceed the geodesic length; we get a condition on wavelength:
$\lambda_{\gamma}\leavevmode\nobreak\ (absorp)\leq 2\pi R_{s}=\lambda_{sm}$ (27)
which is equivalent to the energy quantum absorption condition:
$E_{\gamma}(absorb)\geq\frac{2\pi\hbar c}{\lambda_{sm}}=\frac{\hbar c}{R_{s}}.$ (28)
If, during the specified time interval $\Delta t_{s}$ (22), which corresponds to the duration of the quantum transition from energy level $n_{s}$ to $n_{s-1}$, the black hole does not absorb the aforementioned free light quantum of energy, it will continue to evaporate. If it receives, it grows in accordance with the law:
$\Delta R_{s\gamma}=\frac{2G}{c^{4}}E_{\gamma}\leavevmode\nobreak\ (absorb)\geq\frac{2G\hbar}{c^{3}R_{s}}=\frac{l_{p}}{\sqrt{n_{s}}}$ (29)
while the absorption area $A_{s}$ increases by an amount equal to:
$\Delta A_{s\gamma}\geq 16\pi l_{p}^{2}.$ (30)
It is noteworthy that analogous results prompted Bekenstein to formulate his renowned equation for black hole entropy [51]. However, employing Eqs. (6), (25), and (28) in the context of cosmology, particularly regarding dark matter production, yields the most interesting findings:
$E_{\gamma}(absorb)\geq 2E_{\gamma}(rad)$ (31)
that can be expressed in terms of temperature in the following striking way:
$T_{\gamma}(absorb)\geq 2T_{s}.$ (32)
It means when the temperature of the universe ($T_{\gamma}$) falls below twice the temperature of a black hole, it commences evaporation. Furthermore, it is essential to comprehend how, in the nascent universe, the demise of one generation of PBHs facilitated the emergence of the subsequent generation [14, 15, 58, 59]. We must adhere to these paramount cosmological criteria in our explanation of the formation of contemporary dark matter, which comprises black holes of asteroid-like mass [60]. Another significant consequence of these interactions is that an evaporating black hole can transfer its electromagnetic radiation energy to another black hole only if the mass of the latter is double that of the former [61],
$M_{s_{2}}(received)\geq 2M_{s_{1}}(emitted).$ (33)
These correlations elucidate how, in the primordial phases of the universe’s growth, there occurred a cyclic transfer of energy from a dying, more abundant yet lighter generation of black holes to a new, somewhat less abundant but heavier generation. This is essential for comprehending both the primordial formation of dark matter’s structure and the initial “clumping” of the universe, as these consistent cycles of dark matter generation coincided with the hyper-exponential proliferation of future SMBHs (SMBHs), which served as the centers for the formation of subsequent galaxies.
VII Quantum Information Conservation and Gravitational Radiation from Black Hole collisions
We find that two-particle emission outside the black hole and accompanied by gravitational radiation can potentially resolve the ‘information loss paradox.’ John Preskill [18] and Leonard Susskind [62] define each emitted particle as a carrier of one bit of quantum information, in accordance with the principles of Shannon entropy [63]. This information is represented in two possible spin states, $|0>$ and $|1>$, as per the notation established by Paul Dirac [64]. This mechanism helps keep information safe and shows that the unitarity principle in quantum theory cannot be broken. This means that no information is lost during the process of black holes evaporating.
According to the conservation information law for a binary black hole system ($BH_{1}$,$BH_{2}$), which is derived from the basic unitarity principle of quantum theory, we can write:
$I_{s1,2}=I_{s1}+I_{s2}=2n_{s1}+2n_{s2}.$ (34)
We derive an interesting formula for the energy of gravitational radiation resulting from the collision and merging of a binary black hole system,
$E_{s1,2}(gr(rad))$
$=E_{p}(\sqrt{n_{s1}}+\sqrt{n_{s2}}-\sqrt{n_{s1}+n_{s2}}).$ (35)
This formulation, expressed in terms of quantum numbers, succinctly corroborates Hawking’s esteemed area theorem [65] concerning colliding and merging black holes. During the merger, the resultant black hole experiences a ”ringdown” phase, during which it emits gravitational waves while stabilizing into a final state. Similar to a struck bell that resonates and then diminishes into quiet, the black hole emits energy as it reaches equilibrium—thus the evocative name “ringdown.”
Examining this phase enables scientists to evaluate predictions of Einstein’s general relativity and to validate ideas such as the area theorem, which posits that the overall surface area of a black hole’s event horizon cannot diminish over time. Recently, a group examining data from LIGO’s observations of the ringdown phase presented persuasive evidence [66] supporting the validity of Hawking’s seminal finding.
Now, we obtain the general expression for total gravitational radiation energy for N consecutive collisions and mergers of N black holes as :
$E_{sn}(gr(rad))=E_{p}\Big{(}\sum_{i=1}^{N}\sqrt{n_{si}}-\sqrt{\sum_{i=1}^{N}n_{si}}\Big{)}$
$=\Big{(}\sum_{i=1}^{N}M_{si}-\sqrt{\sum_{i=1}^{N}M_{si}^{2}}\Big{)}c^{2}=\Big{(}\sum_{i=1}^{N}M_{si}-M_{sn}\Big{)}c^{2},$
where
$M_{sn}=\sqrt{\sum_{i=1}^{N}M_{si}^{2}}$ (37)
is the mass of the resulting black hole. The above Eq. (36) with Eq. 37) is very important in understanding and solving the key problem of cosmology:
  1. 1.
    When the mass of a SMBH at the center of a galaxy greatly exceeds the asteroid-like mass of PBHs situated in the SMBH’s “corona,” the dynamics of their mergers yield an interesting outcome. In these instances, nearly all the mass and energy of the tiny PBHs is converted into gravitational wave energy during the merger phase. This process demonstrates a significant transformation in which dark matter is permanently converted into dark energy within the universe. This process serves as a substantial validation of P. W. Anderson’s conjecture, highlighting the intricate interactions among different forms of matter and energy in the universe [67]. Moreover, as will be shown subsequently, these processes account for the universe’s present phase of accelerated expansion.
  2. 2.
    In contrast, Eqs. (36) and (37) suggest that the formation of a SMBH is unlikely during multiple collisions and mergers of black holes with nearly equal masses, as most of their total energy would be dissipated as gravitational radiation. It can be inferred that SMBHs primarily originated from the significant light energy associated with radiative processes in the early universe. We will explain below that these critical insights are vital for precisely characterising and comprehending the substantial role of PBHs and their gravitational radiation in cosmogenesis.
VIII Planck-scale physics, the hot Big Bang, and the Universe’s accelerating expansion
The most astonishing aspect of the proposed technique is the logical and inherent emergence of the Planck unit system, devoid of any extraneous or invalid assumptions and hypotheses. Furthermore, from this unforeseen emergence, we discern the authentic principles, laws, and interactions that ruled the observable universe in its primordial condition. This directly facilitates access to the unexplored realm of Planck-scale physics at the inception of time.
If we traverse backward through the ’arrow of time’, we found spacetime,which is fundamentally three-dimensional and granular; interactions within this framework may lead to photons arranging themselves into a two-dimensional condensate as a lower-energy or stable form. This implies that the 3D spacetime ”background” dictates the conditions that allow the 2D condensate to form and persist. These condensed photons having substantial Plank energy and mass densities, are contained within a distinct, non-Euclidean spacetime, characterized by a circumference-to-radius ratio of one. This unique geometry emerges from the substantial Planck gravitational potential, $\phi_{p}$, which is equivalent to $c^{2}$, thus confirming the peculiar structure of this domain. Notwithstanding its granular and highly energetic characteristics, this 3D Planck photon condensate establishes an exceptionally dense, spherical arrangement. In a broader perspective, this dense condensate functions as a singular entity with a conventional Euclidean metric, integrating the non-Euclidean characteristics of individual photons into a unified, macroscopic structure.
The Planck-Heisenberg uncertainty relation, $E_{p}t_{p}=\hbar$, limits the duration of the observable universe’s very unstable starting state to Planck time, $t_{p}$. Given this tremendous instability, this primordial state soon “explodes,” which is consistent with the predictions of the hot Big Bang theory. This raises several fundamental concerns, like what happens during this explosive event, how long it lasts, and what alterations occur immediately following the Big Bang.
In this first stage, we take the universe as a “granular” spacetime structure that is compact, uniform, and “granular” [68, 69, 70, 71]. Each unit is made up of Planck-scale photons that are packed together into spherical, high-energy packets. Each Planck-condensed photon in this tightly packed, granular spacetime has twelve other photons surrounding it in a very symmetric arrangement. As an example, we can think of the FCC structure of atoms/particles, which is the most dense configuration having twelve nearest neighbors in 3D.
When the initial state is “exploding,” it is critical to evaluate the possibility of several outcomes for these densely packed photons. For example, what are the chances that a photon with twelve neighbors will break out from its packed configuration? Similarly, how likely is it that groups of twelve surrounding photons will form stable Planck-scale binary systems, either among themselves or with adjacent photons?
To quantify these possibilities, one would need to solve a twelfth-degree equation that captures the probability interactions between each photon and its twelve nearest neighbors [4]:
$\aleph=(1-\aleph)^{12}.$ (38)
This equation serves to describe the likelihood of certain structural transformations occurring in the aftermath of the explosion, such as the production of photon pairs, chains, or clusters, which might subsequently spread and contribute to the emergent structure of spacetime and matter in the early universe. Solving this twelfth-degree equation would provide a statistical foundation for understanding the complicated dynamics driving the early microstates and their ability to create larger structures or liberate photons in the immediate aftermath of the Big Bang.
By solving the above Eq. (38) we get the desired probability as: $\aleph=0.1474492855...\approx 0.14745$.
Our focus here is on Planck binaries, as their mergers during the Big Bang would give rise to primordial Planck-scale black holes accompanied by Planck-scale gravitational waves. By the energy conservation law, each Planck binary system, comprising two units of Planck mass-energy, would necessarily convert into the mass-energy of a Planck black hole. Additionally, the work involved in the merger process would inevitably result in the emission of Planck-scale gravitational radiation. Mathematically, the entire process is stated as follows:
  1. 1.
    With probability $\aleph$, condensed Planck photons convert to free ones:
    $(\gamma_{p})\to\gamma_{p}\leavevmode\nobreak\ (free),$ (39)
    the average energy density of free Plank light is
    $\Omega_{L}=\aleph=0.14745.$ (40)
  2. 2.
    Planck binaries ($\gamma_{p},\gamma_{p}$) are transformed into Planck black holes ($BH_{p}$) with probability $-\aleph$, followed by Planck gravitational waves [4]:
    $(\gamma_{p},\gamma_{p})=BH_{p}+\Gamma_{p}$ (41)
    where the energy of gravitational radiation ($\Gamma_{p}$) is caused by Planck work $W_{p}$.
    $E(\Gamma_{p})=W_{p}=F_{p}l_{p}=G\frac{M_{p}M_{p}}{l_{p}^{2}}l_{p}=E_{p};$
    $E(BH_{p})=E(\gamma_{p},\gamma_{p})-E(\Gamma_{p})=2E_{p}-E_{p}=E_{p}.$
    Now we calculate the average Planck dark energy density as
    $\Omega_{DE}=\frac{1-\aleph}{2}\approx 0.42628$ (43)
    The average mass/energy density of Planck dark matter is:
    $\Omega_{DM}=\frac{1-\aleph}{2}\approx 0.42628$ (44)
Take note that current observations indicate that the universe is dominated by dark components, including dark energy and dark matter, which influence its structure and evolution in fundamental ways. Within this framework, it is plausible to consider that the remaining Planck energy density (other than free Planck light energy density) could be statistically divided equally into two distinct contributions: the Planck dark energy density and the Planck dark matter density in its initial stage of formation. This division would suggest that both dark energy and dark matter possess comparable Planck-scale energy densities, potentially offering insights into the underlying nature of these mysterious components that shape the universe. Another aspect is that, by evenly distributing bound energy, we assume that Planck-scale vacuum fluctuations give rise to both dark matter and dark energy states as dual facets of a quantum vacuum, with each contributing equally to the universe’s energy content.
So, we found that the Planck energy densities of fundamental free light, dark matter, and dark energy are spread in the following proportion:
$\Omega_{L}:\Omega_{DM}:\Omega_{DE}\approx 0.14745:0.42628:0.42628$ (45)
Note that if we consider the primary 3D Planck photon condensate to be the “dark” primordial matter, the hot Big Bang process closely resembles the biblical tale of Genesis: “…and the Creator separated Light from Darkness, and there was the First Day ($\aleph$)…” [72]. In the Schroeder [72] book, he suggests that the description of light’s emergence in Genesis aligns with the sudden expansion of light and matter in the Big Bang, paralleling the creation of “light” and separation of “darkness” mentioned in the biblical account. Physically, $\aleph$ delineates a universal scale devoid of any extraneous assumptions, reflecting the stability or instability of particle interactions under high-energy conditions around the Planck scale. It elucidates how PBHs and gravitational waves, together with other remnants from this epoch, may contribute to the Lambda-CDM model as large-scale structures arise from Planck-scale physics. $\aleph$ also plays a significant role in understanding early universe dynamics, marking the likely “separation” behavior of photons and the proportion of primordial energy that remains unbound—eventually contributing to the observable structure of dark matter and dark energy in our current universe.
The so-called “coincidence problem” may have been resolved in a surprising way at the very beginning of time. The breakdown of extreme homogeneity, which separated the initial “graininess” of spacetime into Planck-scale binaries and isolated condensed Planck photons, was caused by spontaneous Planck fluctuations. Planck “granules” acted as “binary explosives,” while Planck fluctuations functioned as “fuses.”
It is also possible to compute the duration of the process in two different methods. The wavelength of a Planck-free photon is $2\pi$ times longer than that of a Planck condensed photon, which is equivalent to the Planck length. Naturally, the implementation of such a wavelength necessitates a duration equivalent to the $2\pi t_{p}$. The second analogous solution arises from the epoch of the formation of Planck black holes. According to the principle of causality, it is equivalent to half the length of its geodesic line divided by the speed of light, specifically $2\pi t_{p}$. This apparent alignment of results is likely, not coincidental; rather, it reflects a profound synchronization inherent to all processes at the universe’s inception, indicative of the fundamental symmetries and coherence present at the time of the Big Bang.
The cosmological constant ($\Lambda$) is a primary parameter in explaining the universe’s accelerated expansion. Calculations indicate that $\Lambda$ could theoretically reach a maximum value around $\frac{1}{l_{p}^{2}}$, where $l_{p}$ is the Planck length. This theoretical peak, about $10^{122}$ times greater than today’s observed value, would correspond to early-universe conditions of intense energy density and rapid expansion. This highest value of $\Lambda$ matches a Hubble parameter on the order of $\frac{1}{t_{p}}$, which is the Planck time. This means that the universe is expanding as quickly as it can under the limits of quantum gravity [73, 74, 75, 76].
The previous “vacuum energy” problem needs to be fixed, which said that the cosmological constant was about $10^{120}$ times bigger than what was observed. This is what has been called the “cosmological constant problem” or “vacuum catastrophe,” which is one of the biggest differences between theory and observation in physics [77, 78, 79].
By proposing that $\Lambda$ reaches its maximum value only under early-universe conditions (near the Planck scale), this approach may nullify the enormous factor of $10^{120}$ and indicates that the observed, much smaller value of $\Lambda$ reflects a natural evolution from this maximum rather than a fine-tuning issue. This revised perspective potentially resolves the cosmological constant problem by aligning $\Lambda$ with realistic early-universe scenarios, where its value would have been vastly higher than it is now.
Now the question is: what happened during the subsequent “days of creation”? Subsequent cosmic events start to unfold in full accordance with the principles outlined in the preceding sections.
The predominant portion of energy from the Planckian free light is absorbed by a fraction, $\aleph$ (40), by the Planck black holes, effectively doubling their mass/energy density. This absorption leads to a reduction in the dark matter mass/energy density to $\Omega_{DM}=2\aleph$. Simultaneously, the residual segment of the Planck black holes undergoes disintegration, yielding two photons $\gamma_{p,0}$ and a gravitational wave $\Gamma_{p,0}$:
  1. 1.
    With a probability $\aleph$ the next generation of black holes is formed by the following scheme:
    $\gamma_{p}+BH_{p}\to BH_{s}\leavevmode\nobreak\ (n_{s}=4)$ (46)
    Here, to calculate $n_{s}$, we have to use Eqs. (12) and (42), and correspondingly
    $M_{s}=M_{p}\sqrt{4}=2M_{p},$ (47)
    and average (typical) dark matter mass/energy density
    $\Omega_{DM}=2\aleph\approx 0.29489.$ (48)
  2. 2.
    With a probability $\frac{1-\aleph}{2}-\aleph=\frac{1-3\aleph}{2}$ the plank blackhole decay occures:
    $BH_{p}\to 2\gamma_{p,0}+\Gamma_{p,0}$ (49)
    with
    $2E(\gamma_{p,0})=2\times\frac{E_{p}}{4},$
    $E(\Gamma_{p,0})=\frac{E_{p}}{2}$ (50)
    where we have used Eqs. (15) and (49), and average incremented dark energy:
    $\Delta\Omega_{DE}=\frac{1-3\aleph}{2}\times\frac{1}{2}=0.13941,$ (51)
    and the average irreversible dark energy density:
    $\Omega_{DE}=\Omega_{DE}\leavevmode\nobreak\ (precedent\leavevmode\nobreak\ phase)+\Delta\Omega_{DE}$
    $=\frac{1-\aleph}{2}+\frac{1-3\aleph}{4}=\frac{3-5\aleph}{4}\approx 0.56569$ (52)
    and the corresponding average energy $\Omega_{L}$, emitted through the Planck $BH_{p}$ decay to $\gamma_{p,0}$ photons is
    $\Omega_{L}=1-\Omega_{DM}-\Omega_{DE}=\frac{1-3\aleph}{4}\approx 0.13941.$ (53)
    However, by the direct calculation, we have the same result as follows:
    $\Omega_{L}=\frac{1-3\aleph}{2}\times 2\times\frac{1}{4}=\frac{1-3\aleph}{4}\approx 0.13941.$ (54)
When we compare this second phase (“the second day of creation”) with the preceding (Big Bang) one, we find that the dark energy density at Big Bang increased from zero to $0.42628\leavevmode\nobreak\ (=0.56569-0.13941)$, however, during the second part of the Planck epoch, it only grows with the value of $\Delta\Omega_{DE}\approx 0.13941$. This implies a significant drop in accelerated expansion in the “baby universe” (when this increment stops, the universe enters the stage of non-accelerated expansion). We must note that, in line with this process, the density of dark matter decreases from $0.42628$ (44) to $0.29489$ (48).
From this point onward, the gravitational mass of the universe, represented solely by dark matter, will continually decrease, acting as an inexhaustible reservoir of energy for all future generations, leading to the formation of contemporary dark matter, as well as the subsequent emergence of extremely massive and SMBHs and “ordinary” matter. Let us continue with the examination of the third phase (the “third day of creation”) of the Planckian era. By quantum absorption, the formation of typical black holes $BH_{s}$ ($n_{s}$=4) will quickly grow to the standard creation of $BH_{s}$ ($n_{s}$=8).
Now the question is: What is the basis for this knowledge? According to the preceding phase, the maximum energy reserve for the creation of a typical primordial black hole $BH_{s}$ ($n_{s}$=4) via the quantum absorption of free light must be articulated as the total of the prior phase:
$\Omega_{DM}(max)=\Omega_{L}+\Omega_{DM}=1-\Omega_{DE}=\frac{1+5\aleph}{4}$
where we have used Eq. (52). In terms of a typical black hole, this gives the average mass
$M_{s}=M_{p}\frac{\Omega_{L}+\Omega_{DM}}{\aleph}=\frac{1+5\aleph}{4\aleph}\approx 2.945498M_{p}$ (56)
Nonetheless, the value 2.945498 does not belong to the series represented by $\sqrt{n_{s}}$, as it is required to be a conventional quantum number. According to the principle of maximum black hole entropy $S_{s}$, we determine that the maximum value of $n_{s}$ that meets the natural requirement is $\sqrt{n_{s}}<2.945498$. As a result, we get $n_{s}$=8, the required number for a typical third-generation primordial black hole.
Using previously obtained results, we find the lifetime (using Eq. (23)):
$\tau_{s}=\sum_{n=1}^{n_{s}=8}2\pi t_{p}\sqrt{n}\approx 32.6\pi t_{p},$ (57)
and emitted free light density (using Eq. (15)):
$\Omega_{L}=\aleph\sum_{n=1}^{n_{s}=8}\frac{1}{2\sqrt{n}}\approx 0.32228.$ (58)
This is immediately transformed into the $\Omega_{DM}$ density of the subsequent generation of PBHs via radiative transfer.
The initial findings suggest that taking into account the earlier stages, the total length of the Planckian era is bounded by the value $50\pi t_{p}\approx 8.46\times 10^{-42}$ seconds.
The second result indicates the complete extinction of PBHs that emitted significant light energy, thereby enabling the formation of the next generation. Prior research concerning the principles of quantum absorption indicates that, in the context of radiative transfer between black holes, the mass of the latter must be not less than twice that of the former.
However, what is the source of the final entities? To this point, our analysis has been confined to average, representative values concerning the formation of black holes. In accordance with the foundational concepts of probability theory, it is essential to examine a comparatively small yet impactful quantity of black holes capable of producing a new generation characterized by mass/energy density,
$\Omega_{DM}\approx\Omega_{DL}\approx 0.32228.$ (59)
As nearly all light energy was absorbed by this generation of PBHs, we arrive at a significant conclusion: by the close of the Planckian era, dark energy approaches its maximum value,
$\Omega_{DE}\approx 1-\Omega_{DM}\approx 0.67771$ (60)
and its final increment:
$\Delta\Omega_{DE}\approx 0.67771-0.56569\approx 0.11202.$ (61)
The latter indicates that the conclusion of the Planckian era marks the termination of the initial phase of the universe’s accelerated expansion, transitioning to a phase characterized by non-accelerated nevertheless very rapid expansion, as demonstrated below.
Comparing the derived dark energy density value of $0.6777$ with the recent observational data of $0.6847$ from the Planck Collaboration (2020) [80] suggests that the universe has experienced accelerated expansion over the past five to six billion years due to this increment in dark energy density. This change is given by
$\Delta\Omega_{DE}=\Delta\Omega_{\Lambda}\geq 0.6847-0.6777$
$=0.0073\leavevmode\nobreak\ (or\leavevmode\nobreak\ greater)$ (62)
and varies according to differences observed in recent results from various cosmological measurement methods [80, 81]. It is also noteworthy that primary gravitational waves (radiation), ($\Gamma_{p}$ and $\Gamma_{p,0}$), with energies $E_{p}$ and $\frac{E_{p}}{2}$ respectively, account for approximately $97\%$ of the total dark energy density in the present-day universe, aligning with the findings from recent observational studies [80].
IX The growth of supermassive black holes and very early clumping of the universe
This section addresses the challenge of understanding the formation mechanisms of SMBHs and the initial clumping phenomena observed in the early universe.
As established earlier, successive generations of black holes formed primarily through the radiative transfer of light energy. This process involved energy transfer from one generation, which was significantly more numerous but “lighter,” to a subsequent generation that was far less numerous but “heavier.” This transfer of energy allowed for the growth and evolution of PBHs over time, leading to the formation of SMBHs at the centers of galaxies. These SMBHs play a crucial role in shaping the structure and dynamics of galaxies. This can be articulated straightforwardly and lucidly with significant implications,
$\mathcal{N}_{sj}E_{p}\sqrt{n_{sj}}=\mathcal{N}_{sj-1}E_{p}\sqrt{n_{sj-1}}$
$=\Omega_{DM}E_{u}=0.32228\times 1.5\times 10^{62}E_{p},$ (63)
where $\mathcal{N}_{sj}$ and $\mathcal{N}_{sj-1}$ represent the average numbers of the PBHs in a given and previous generations respectively, while $n_{sj}$ and $n_{sj-1}$ refer to the typical quantum numbers representing these generations, $E_{u}$ is the total energy of the observable universe. It is easy to understand that the time gap between generations corresponds to the lifespan of the preceding generations. According to Eq.(23) we can write,
$\Delta t_{sj,sj-1}=\frac{4\pi}{3}t_{p}(\sqrt{n_{sj-1}})^{3}.$ (64)
Hence, it is easy to observe a direct relationship between the $\sqrt{n_{s}}$ and the time as:
$\sqrt{n_{s}}\propto t^{\frac{1}{3}}.$ (65)
Taking into account that the cosmological scale factor $a(t)$ is proportional to $\sqrt{n_{s}}$, we find that,
$a(t)\propto\sqrt{n_{s}}\propto t^{\frac{1}{3}},$ (66)
so that the Hubble parameter is
$H(t)=\frac{\dot{a}}{a}=\frac{dln(t^{\frac{1}{3}})}{dt}=\frac{1}{3t}$ (67)
and Hubble radius
$R_{H}(t)=\frac{c}{H(t)}=3ct.$ (68)
Our selection of the scale factor $a(t)$ as proportional to $\sqrt{n_{s}}$ is substantiated by the ideas of Bekenstein and Hawking [14, 51], which assert that entropy is intrinsically associated with the surface area of a black hole. In our proposed model, this surface area is proportional to $n_{s}$. Therefore, it may be inferred that we should consider taking scale factor a to be proportional to $\sqrt{n_{s}}$. Alternatively, we can explain the reason as following: if we consider $a(t)\propto\sqrt{n_{s}}$, this proportionality allows a simplified relation where changes in $a(t)$ directly reflect variations in the number density in the early universe, to simplify the description of the universe’s expansion concerning matter or radiation densities [83, 84, 85].
Now we can verify our findings by using the fact that the universe’s age and observable radius are [83, 86]: $Age_{U}\approx 8.08\times 10^{60}t_{p}$ and $R_{U}\approx 2.7\times 10^{61}l_{p}$
We find the Hubble Radius at the Universe’s age
$R_{H}(Age_{U})=3c\times 8.08\times 10^{60}t_{p}=2.424\times 10^{61}l_{p}$ (69)
Two significant insights can be gained from this unexpected consistency of our results: first, that the suggested method effectively addresses the challenging cosmological problems; and second, that it validates the rapid, non-accelerated expansion of the very early universe ($q=-\frac{\ddot{a}a}{\dot{a}}=2>0$).
The difference of $-10\%$ between the calculated Hubble radius ($R_{H}(Age_{U})$) and the radius ($R_{U}$) of the present universe clearly shows that the latter has grown a lot because of the second phase of faster expansion. Having comprehended the evolution of dark matter throughout generations, we are prepared to address the puzzling issue of the origin and growth of SMBHs.
Immediately after the end of the Planckian era, an extraordinary event began to unfold: a few exceptionally rare PBHs gained masses far exceeding those of typical black holes. Each death of the previous generation released an immense burst of light energy, with a small fraction absorbed by the rapidly growing accretion layers of these future cosmic giants.
Based on the fact that both the quantum information and the growing black hole’s accretion area are proportional to the quantum number, we can write a growth equation for $n_{s}(t)$ in accordance with Bohr’s correspondence principle as:
$\frac{dn_{s}(t)}{dt}=B_{s}(t)n_{s}(t),$ (70)
where $B_{s}(t)$ is the accretion (light absorption) function, expressing the dependence on the radiation of each generation of typical black holes evaporation. We find that
[…]
As a result, we obtain (using Eq. (6)) the exponential growth of the mass of a supermassive black hole:
[…]
It should be noted that the time-dependent mass function can also be found in dynamical spacetimes like those of the Vaidya spacetime [87, 88, 89], and aligns with models involving time-dependent mass functions for SMBHs [90].
The growth of it’s radius is determined in the same way (Eq.(6)) as:
[…]
The growth of SMBHs occurs through a cyclical mechanism in which successive generations of PBHs generate significant photon emissions that facilitate the growth of subsequent black holes. Fundamental Eqs. (15) and (24) describe this cyclic mechanism, which results in significant increases in luminosity with each cycle while the SMBH’s accretion area expands at an even faster rate, enhancing its gravitational pull. The gravitational influence promotes the aggregation of PBHs, which signify the evolution of dark matter in the early universe. The duration (23) of each subsequent cycle increases significantly, thereby improving the formation efficiency of early galactic precursors and SMBHs of different sizes. As these cycles end, the universe cools, leading to a significant reduction in the growth of SMBHs and signaling a shift to a less active cosmological phase.
[…]
The rapid expansion of SMBHs contributes to early cosmic clustering, which aids in the formation of protogalaxies and, subsequently, early galaxies. Observations from the James Webb Space Telescope (JWST) support this scenario, demonstrating that massive and supermassive black holes act as gravitational attractors, facilitating the formation of large cosmic structures in the early universe.
X Baryogenesis through primordial black hole production
The most convincing evidence of our model is that PBHs are directly tied to baryogenesis. It is well-known that dark matter surrounds baryonic matter everywhere [93, 94, 95]. Moreover, we find, in general, that where the concentration of dark matter is higher on large scales, there is a corresponding increase in the concentration of ordinary matter.
Strictly speaking, the two-particle emission process as represented in Eq. (11) should be written in the following general form:
$BH_{s}\to BH_{s-1}+X_{s}+\bar{X}_{s}+\Gamma_{s}$ (74)
where $X_{s}$ and $\bar{X}_{s}$ represents particle and antiparticle, $\Gamma_{s}$ represents gravitational wave. Since photons are their own antiparticles, we can replace this pair with term $2\gamma_{s}$ and obtain previously presented Eq. (11).
But, when dealing with the production and emission of various particle-antiparticle pairs, the substitution of two $\gamma_{s}$ particles remains acceptable due to their inevitable annihilation. The only exception is when CP violations occur owing to certain gravitational influences.
The unresolved puzzles surrounding baryon asymmetry, baryogenesis, and leptogenesis persist as fundamental challenges within the realms of cosmology and particle physics [96, 97]. The established theories of baryogenesis and leptogenesis exhibit shortcomings in two fundamental aspects. First, they fail to sufficiently consider the observed prevalence of baryonic matter, which constitutes roughly $5\%$ of the total mass-energy of the universe [80]. Second, they fail to fully explain the precise balance between proton and electron numbers, a symmetry that yields net charge neutrality at both microscopic and macroscopic levels across the “Cosmic Web” [98].
Instead, we will take a different approach—moving from facts to theory—by looking at the unique productivity of PBHs and their distinct gravitational and thermodynamic effects.
We propose to treat the observed baryon asymmetry as a fundamental postulate in the same way that Einstein accepted the constancy of the speed of light based on empirical evidence. By taking this asymmetry as a given, we can then explore its implications within the broader framework of cosmological and particle-physics phenomena.
Our focus lies in establishing a correlation between this imbalance and the mechanisms underlying particle pair formation by black holes, utilizing a systematic approach. This framework makes it possible to find strong links between asymmetry in the real world and mechanisms that work in theory. This helps us learn more about how particles form in the early universe and when gravity is very strong.
In the high-energy environment of the early universe, neutrons and protons could freely convert into one another due to the abundant energy available [22, 99]. However, as the universe expanded and cooled, the energy levels dropped, particularly when positron-electron ($e^{+}-e^{-}$) annihilation became significant and neutron decay to protons became more prevalent, as lower available energy could no longer support reverse conversions [85]. In this context, PBHs start to play a role in contributing to ordinary matter through decay processes [11, 23, 61]. The decay interactions of PBHs can be represented as follows:
$BH_{s}\to BH_{s-1}+2n^{0}+\Gamma_{s}$ (75)
where each decay step produces two neutrons ($n^{0}$) and an gravitationl energy ($\Gamma_{s}$), associated with the process. Subsequently, the neutrons decay further:
$2n^{0}\to 2p^{+}+2e^{-}+...$ (76)
Here, $n^{0}$, $p^{+}$ and $e^{-}$ represent neutrons, protons, and electrons respectively. Through these sequential interactions, PBHs can contribute to the production of baryonic matter, aiding in the synthesis of fundamental particles necessary for nucleosynthesis and the eventual formation of atomic nuclei [16, 73, 83, 99]. Now, we can easily determine the restriction imposed on the “productive” quantum numbers of the respective PBHs following Eq. (25),
$\frac{E_{p}}{4E(c_{q})}\leq\sqrt{n_{s}}\leq\frac{E_{p}}{4E(n^{0})}$ (77)
where $E(n^{0})=0.9396GeV$ is the rest energy of the neutron, $E(c_{q})=1.25GeV$ energy of the free “charm” quark, next in this parameter after neutron. This constraint implies that PBHs predominantly generate neutrons up to a specific quantum energy level $n_{s}$. The highest limit, $E(c_{q})$, ensures that the PBHs won’t have enough energy to produce charm quarks, restricting particle production to neutrons and lighter particles. By limiting $n_{s}$ to the specified range, the model ensures that the PBH decays predominantly yield baryons, such as protons, rather than other particles that would not enhance baryon asymmetry. The concept assumes that gravitational effects or primordial black hole decay can induce CP violation, essential for generating a matter-antimatter imbalance. PBHs create conditions in which gravitational interactions may inherently result in such violations, eliminating the necessity for alternative CP-violating mechanisms. Here we need the “charm” quark rest energy in order to determine the quantum number $n_{s}(c_{q})$ at which neutron production stops. Using a simple equation
$\frac{[2n^{*}_{s}(n^{0})-2n_{s}(c_{q})]E(n^{0}){\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}]}}{E_{p}\sqrt{n_{s}^{*}(n^{0})}}=\frac{\Omega_{b}}{\Omega_{DM}}$ (78)
(where $\Omega_{b}$ is modern baryonic density) we can determine the typical boundary on $n_{s}^{*}(n^{0})$, which characterises the generation of black holes capable of producing “ordinary matter” of the observable universe.
The expression on the left side of Eq. (78) denotes the productivity of one typical black hole, expressed as the difference between the number of neutrons and the number of annihilated charm quarks it produces (square brackets) multiplied by the rest of the energy of the neutron and divided by the rest energy of typical producing black hole. We equate this to the energy density of baryon $\Omega_{b}$=0.0496 divided by the energy density of the producing black hole $\Omega_{DM}$=0.32228 (Eq. (59)). In essence, this equation as a whole expresses a delicate balance between the productive capacity of PBHs in generating baryonic matter and the observed ratio of baryonic to dark matter in the universe.
Here, we find a quadratic equation for $\sqrt{n_{s}^{*}(n^{0})}=x$ as:
$x^{2}-\frac{\Omega_{b}}{\Omega_{DM}}\frac{E_{p}}{2E(n^{0})}x-n_{s}(c_{q})=0$ (79)
Solving the above Eq. (79) by using $\sqrt{n_{s}(c_{q})}=\frac{E_{p}}{4E(c_{q})}$ and the values of $\Omega_{DM}$, $\Omega_{b}$ and $E_{p}=1.22\times 10^{19}$eV, we find that $\sqrt{n_{s}^{*}(n^{0})}=2.9902\times 10^{18}$ corresponding to energy $E^{*}(n^{0})=1.02GeV$. Here, we disregard the negative root of the quadratic equation, as the quantum number $n_{s}$ cannot take negative values. Consequently, ordinary (baryonic) matter production from PBHs is feasible within the energy range of $1.02$ to $1.25$ GeV.
According to Eq. (23) the formation of baryonic matter takes place over roughly $87$ thousand years, while the complete cycle of primordial black hole evaporation and subsequent extinction lasts approximately $190$ thousand years. During the intermediate period, from $87$ to $190$ thousand years, PBHs may continue decaying and emitting radiation, even though baryonic matter production has largely ceased. This radiation increases the universe’s overall radiation density but does not significantly affect the baryon-to-dark matter ratio, which was established during the initial baryonic production phase [102].
To determine the total period during which neutron-to-proton/electron conversion occurs, we need to calculate the duration of proton/electron generation. Based on the above information, we can infer that this conversion process begins when the available energy decreases below the rest mass energy of the charm quark ($1.25$ GeV) and concludes once the available energy reaches the threshold value (i.e., $1.02$ GeV) for baryonic matter formation. For this process, the corresponding quantum numbers are $\sqrt{n_{si}}=\frac{E_{p}}{4E(c_{q})}=2.44\times 10^{18}$ and $\sqrt{n_{sf}}=2.9902\times 10^{18}$ (Eq.(79)), where $n_{si}$ and $n_{sf}$ stand for the initial and final quantum number for the generation of baryonic matter.
This duration corresponds exactly to half of the cosmic interval from the Big Bang to the Recombination epoch, around $378$ thousand years [80, 103]. This epoch also witnessed the final significant reheating event in the universe. As the universe cooled, ordinary matter settled into a relatively hot plasma state, enabling primary nucleosynthesis, during which protons and neutrons combined to form the first atomic nuclei, signaling the beginning of a phase in which stable elements could emerge.
The most important observation is that the processes involved in the generation of baryonic matter, the evaporation of black holes, and the reheating phase occurred in localized “clumps” rather than in a uniform manner. This non-uniformity played a significant role in the subsequent formation of the first galaxies, suggesting that before the phase of ordinary matter creation, the universe likely displayed inhomogeneities. The quantum perturbation [22] could also give rise to these inhomogeneities. Primordial density fluctuations can lead to the formation of gravitational wells that attract matter, thereby increasing particle interactions, promoting the formation of black holes, and resulting in the production of matter within those regions. Consequently, primordial galaxies emerged in proximity to these dense regions, where primordial black holes facilitated baryonic synthesis and localized reheating occurred, thus establishing the foundation for the large-scale structure of the Universe.
XI Formation of Modern Dark Matter and Large-Scale Cosmic Structure
The mechanisms of baryogenesis and leptogenesis, along with the continuous radiative transfer of light energy from the last generation of evaporated PBHs to their contemporary equivalents, persist indefinitely.
The formation process ended when the temperature inside the dense clumps dropped below the threshold level of $511\leavevmode\nobreak\ KeV/k_{B}$, which is the electron rest energy, and then cooled to a lower equilibrium value of $80\leavevmode\nobreak\ KeV/k_{B}$ [104]. At this stage, the opacity within the medium was significantly reduced, stopping the growth of PBH, which is posited as the beginnings of dark matter [23]. Consequently, these PBHs evaporated through Hawking radiation, emitting high-energy X-rays into the surrounding environment.
By applying the Eqs. (6), (11) and (15) that relate the threshold accretion temperature to the mass of a black hole and its associated parameters, we derive the following typical characteristics of contemporary dark matter:
$\sqrt{n_{s}}=\frac{E_{p}}{2\times 80KeV}=7.625\times 10^{22}\leavevmode\nobreak\ (typical),$
$M_{s}=1.659\times 10^{18}g\leavevmode\nobreak\ (typical);$ (80)
$\sqrt{n_{s}}=\frac{E_{p}}{2\times 511KeV}=1.194\times 10^{22}\leavevmode\nobreak\ (minimal)$
$M_{s}=2.598\times 10^{17}g\leavevmode\nobreak\ (minimal),$ (81)
where we have to use Planck mass $M_{p}=2.176\times 10^{-5}g$.
The hard X-ray spectrum emitted by PBHs offers valuable insights into the nature of dark matter and the mechanisms potentially driving cosmic acceleration. According to Eqs. (31), (80), (81) this emission spectrum is characterized by a peak intensity at photon energies near $E_{\gamma}=30\leavevmode\nobreak\ KeV$ and typical intensity of $E_{\gamma}=40\leavevmode\nobreak\ KeV$, with an upper cutoff observed at around $250\leavevmode\nobreak\ KeV$. This exactly matches the unresolved origin of the cosmic hard X-ray background (CXR), which was discovered before the famous cosmic microwave background (CMB). The spectral properties are important as they may clarify the mass distribution of PBHs and impose constraints on dark matter theories, considering the proposed role of PBHs as candidates for dark matter [11, 23]. Furthermore, the X-ray emissions contribute to our understanding of high-energy processes in the early universe, influencing theories on the acceleration of cosmic expansion.
This distinctive hard X-ray spectrum has been observed in high-energy environments across the cosmos, including active galactic centers, galaxy clusters, and superclusters [105]. The extensive emissions indicate the potential existence of a substantial population of PBHs, supporting the theory that PBHs may represent a considerable fraction of dark matter [11, 106]. PBHs may crash into and merge in the hot, high-energy parts of SMBH coronas because of the strong gravitational fields around them. This can release a lot of X-rays. The resulting intense X-ray radiation, especially in non-jet-related phenomena, aligns with the theory that PBH activity may account for a considerable portion of the unresolved cosmic X-ray background [100, 107].
Thus, we may conclude that the hard X-ray spectrum produced by PBHs offers essential information regarding the characteristics of dark matter, and the mechanisms responsible for X-ray emission in regions of high density, which also play a role in the generation of dark energy. These interactions enhance our comprehension of PBHs as candidates for dark matter and indicate their potential importance in the universe’s accelerating expansion, illustrating the relationship between dark matter and dark energy in cosmic evolution. The transition of dark matter into dark energy may possibly resolve the “core-cusp problem” [108].
The JWST has revealed unexpectedly mature galaxies in the early universe, which may be explained by the rapid initial clustering of dark matter halos. This clustering generated significant gravitational potentials that drew ordinary matter into dense regions, accelerating early galaxy formation [109]. The fast growth of these dark matter halos, which create gravitational wells, would naturally speed up the formation of the universe’s structure by directing normal matter into areas with lots of it, where galaxies and stars may come together faster than was thought before. Consequently, these well-structured galaxies, identified in what was previously considered a primordial and underdeveloped epoch, exemplify the dynamic efficacy of the universe’s formative mechanisms.
To the best of our knowledge we can say that black holes with masses around $10^{18}$ g (Eq.(80) have lifetimes comparable to the age of the Universe. Heavier black holes can be considered stable, while lighter ones are actively evaporating or may have already fully evaporated since their formation. Black holes with masses near $10^{18}$ g, however, could still persist today and potentially contribute to dark matter. It is clear from the above statement that the average number density of dark matter particles (or PBHs) is about $10^{41}$ times lower than that of normal matter when we look at the current mass parameter $M_{s}(\sim 10^{18}g)$ along with the known values of $\Omega_{b}$ and $\Omega_{DM}$. This is shown by the formula $\frac{\Omega_{DM}m_{p}}{\Omega_{b}M_{s}}$, where $m_{p}=1.67\times 10^{-24}g$ is the mass of a proton. This aligns perfectly with the cosmic phenomenon of the “Bullet Cluster” [110], where vast clouds of dark matter move through each other without obstruction, demonstrating their minimal interaction. In contrast, intergalactic gas clouds, which possess a notably higher density and are rich in ordinary matter, experience intense electromagnetic interactions that lead to their deceleration and dispersion when they collide. This notable disparity underscores the intriguing properties of dark matter, which interacts exclusively through gravitational forces.
It is important to note that, in the initial epochs of the universe, galaxies were closely clustered, forming a substantial “universal supercluster” filled with matter. As the universe expanded rapidly, fractures and separations emerged, leading to the formation of the massive cosmic structures we observe today—immense voids, superclusters, great filaments, and walls. The important features arise as natural consequences of this rapid expansion, logically derived from the previously mentioned results.
XII Discussion and Conclusion
Our article examines PBHs as crucial factors in the universe’s evolution, specifically the early universe–relating to the Planck scale physics, particularly in relation to quantized gravity and cosmic events. The suggested model links the quantization of black hole characteristics with fundamental physics and cosmology, providing insights into high-energy astrophysics, cosmic structure development. This concept achieves the unification of quantum physics and gravity by the groundbreaking discovery of condensed light, namely, the self-gravitating photon Bose-Einstein condensate. This technique is fundamentally based on a geometric understanding that connects the Compton wavelength of the quantum photon condensate to the gravitational radius of PBHs. This connection elucidates a range of quantized attributes—encompassing geometrical, gravitational, thermodynamic, informational, and radiative properties—illuminating the fundamental principles that dictate the formation, development, and ultimate extinction of successive generations of PBHs.
This advancement unveiled new aspects of Planck-scale and primordial black hole physics, facilitating the creation of efficient and robust mathematical instruments. These tools have facilitated accurate computations and resolutions to various fundamental issues in physics and cosmology, encompassing the beginning of time, Big Bang cosmology, the genesis and development of dark matter and dark energy, as well as the mystery of baryogenesis and baryonic asymmetry. The model also examines the hyper-exponential growth of SMBHs, the birth of primordial galaxies, the evolution of large-scale structures, and the Universe’s two-phase accelerated expansion. Furthermore, it offers very plausible solutions for persistent problems such as the horizon problem, the cosmological constant problem, the information loss conundrum, and the core-cusp dilemma.
The most striking aspect of our findings is that they almost align remarkably well with the predictions and observations of $\Lambda$CDM cosmology. This agreement is particularly notable because our results were derived independently, without relying on the standard equations and assumptions typically associated with $\Lambda$CDM.
It should be noted that, in our article, a method based on the membrane formalism of Price and Thorne [39], through an algebraic technique, is used to show signs of inhomogeneity caused by quantum fluctuations at the Planck scale. These results may fit with recent progress in inhomogeneous cosmology, as pointed out by Sarkar [111], who looked at models that are different from the homogeneous FLRW framework in great detail. Specifically, Sarkar’s work underscores the need for alternative mechanisms of structure formation, particularly in inhomogeneous and anisotropic universes, as discussed by Krasinski [112]. While inhomogeneous models have not yet been as thoroughly explored as FLRW models in terms of structure formation, empirical data increasingly supports a non-uniform cosmic landscape [113, 114, 115, 116]. These observations collectively indicate that large-scale inhomogeneities may play a significant role in the evolution of cosmic structures, suggesting a potential re-evaluation of our standard cosmological models. This article does not address inhomogeneity in detail, as it falls outside the scope of this study.
This research explores the relationship between quantum mechanics and cosmology by analyzing PBHs in terms of their quantum properties and making comparisons with BECs. The founding relationship can be experimentally investigated within the framework of analog gravity, building on the foundational work of Belgiorno et al. [33] and Agullo et al. [34]. The ultimate goal is to develop a unified theory integrating quantum mechanics and general relativity, paving the way for future theoretical advancements and space exploration.
Acknowledgement: The authors extend their gratitude to Prof. R. Ruffini, Prof. G. Vereshchagin, Prof. V. Mukhanov, Prof. S. Ansoldi, Prof. A. Pitteelli, Prof. A. Gaina, Prof. V. Galamaga, Prof. P. Panchadhyayee, I. Borşevici, and S. Borşevici for their valuable and insightful discussions, which greatly contributed to the development of this work.
Conflicts of interest: The authors declare no conflicts of interest.
Data availability: There is no associated data with this article, and as such, no new data was generated or analyzed in support of this research.